# Enhancing the Efficiency of Time-Dependent Density Functional Theory Calculations of Dynamic Response Properties

### v1

@@ 5 Discussion
State-of-the-art XFEL facilities and modern backlighter set-ups enable the investigation of material properties under new, hitherto unexplored conditions 13, 15, 14, 19. While advancements in experimental capabilities have greatly enhanced our understanding of high-pressure physics and laser-matter interaction, they also present new challenges for ab initio simulation techniques. It has become crucial to improve the efficiency of ab initio modeling due to the significant increase in computing power that is required to accurately describe the properties of materials under extreme conditions 27, 86, 91, 92. One key property measured directly in experiments is the XRTS spectrum. Modeling these measurements is a computationally highly intensive task due to strong thermal excitations, the peculiarities of the detectors used to capture the XRTS spectrum, and the variations in temperature and density within the experimental targets. Consequently, developing new computational strategies to improve the efficiency of XRTS modeling has become a crucial task for providing sustainable theoretical support for XRTS experiments, as well as for extending the benchmark and refinement of simplified models commonly used in areas such as radiation hydrodynamics.
@@
In this work, we developed an approach based on the one-to-one mapping between the electronic DSF $S(\mathbf{q},\omega)$ and the ITCF $F(\mathbf{q},\tau)$. This method combines the $\eta$-convergence test in the imaginary time domain with a constraints-based noise attenuation technique. This combined approach constitutes a highly efficient strategy for XRTS modeling, providing a speed-up of up to an order of magnitude and thus potentially saving millions of CPU hours when describing a single XRTS measurement of a target at extreme conditions. The application of the presented workflow was demonstrated on the example of solid-density hydrogen with partially degenerate electrons under conditions relevant for IFE and laser-driven hydrogen jet experiments. A second example is given by isochorically heated aluminum with hot electrons within an fcc lattice structure of ions. This scenario is pertinent to the physics of transient states studied in experiments that utilize X-ray-driven heating 15, 2, 93, 94, 24 (also see discussions in Ref. 65).
@@
By significantly reducing the computational burden associated with modeling XRTS experiments, the presented approach is expected to play a vital role in computational efforts aimed at supporting new experimental initiatives. In addition to XRTS modeling and interpretation, the presented workflow also enhances the efficiency of modeling a plethora of dynamic response properties that are related to the DSF through the fluctuation-dissipation theorem and the Kramers-Kronig relations. Instead of the DSF, the workflows can be initiated using any dynamic response property that can be expressed in terms of the dynamic density response function  51, 95, 96, such as dynamic conductivity from the Kubo-Greenwood method 51, 74, 97, 98 (as outlined in Sec. A.3) or TDDFT 99, 100.
@@ A.1 TDDFT simulation details
The TDDFT simulations were carried out using Quantum ESPRESSO 101, 102, 103, 104, 75, using the Liouville-Lanczos (LL) method to compute the DSF. A static (adiabatic) GGA approximation for the exchange-correlation (XC) kernel was employed in all calculations. For hydrogen, we utilized the pseudopotential h_pbe_v1.4.uspp.F.UPF 105 in conjunction with the PBE XC functional 106. We considered a density value of $\rho=0.08~\text{g/cc}$ and a temperature of $T=4.8~\text{eV}$, corresponding to the Fermi temperature of the electrons 81. The size of the simulation box, which contains 14 protons, is set to $L=6.641174\textup{~\AA }$. This temperature leads to a partially degenerate state, with partial ionization 107, 72, 108, 64. The DSF results were averaged over 20 snapshots, using a Monkhorst-Pack $k$-point grid set to $10\times 10\times 10$. The convergence of the results with respect to the $k$-point grid was verified by testing an individual snapshot with different $k$-point grids: $8\times 8\times 8$, $10\times 10\times 10$, $14\times 14\times 14$, and $18\times 18\times 18$. The snapshots of protons were randomly selected from the particle trajectories generated by the KSDFT-based molecular dynamics simulations. To obtain $10^{5}$ Lanczos coefficients, we employed a bi-constant extrapolation scheme following the computed number of Lanczos iterations $N_{\rm iter}=2\times 10^{4}$ 75. The calculations were executed with a cutoff energy of $75~\text{Ry}$ and $280$ KS orbitals.
For Al, we computed the DSFs in the [100] crystallographic direction and employed the GGA-level PBE functional along with the pseudopotential Al.pbe-sp-van.UPF 109, 110, which accounts for 11 valence electrons, and an energy cutoff of 200 Ry. We employed a cubic simulation box containing the conventional fcc unit cell (4 atoms) with an fcc lattice parameter of $a=4.05\,\textup{~\AA }^{-1}$ 111. At the temperature of $T=0.025\,\text{eV}$, we used a $40\times 40\times 40$ k-point grid. At the higher temperature of $T=25\,\text{eV}$, we employed two different k-point grids: $10\times 10\times 10$ and $18\times 18\times 18$. Additionally, at $T=25\,\text{eV}$, we utilized 400 KS bands for the cell with 4 Al atoms.
@@ A.2 Connection between different dynamic response functions
Here, we briefly outline the relationships between the DSF and the dynamic conductivity as well as the absorption coefficient, which are used to compute the results presented in Section 4.2. For more details, see, e.g., Refs.112, 113, 74. The DSF is connected to the imaginary part of the density response function through the fluctuation-dissipation theorem, as indicated in Eq. (8). This means that the fluctuation-dissipation theorem and the Kramers-Kronig relations allow one to compute the density response function $\chi(\mathbf{q},\omega)$ from the DSF. The dynamic dielectric function can be expressed in terms of $\chi(\mathbf{q},\omega)$ as
@@
The absorption coefficient is computed using the real parts of the dynamic conductivity $\sigma_{1}(\mathbf{q},\omega)$ and dielectric function $\varepsilon_{1}(\mathbf{q},\omega)$ using:
$\alpha(\mathbf{q},\omega)=\frac{4\pi\sigma_{1}(\mathbf{q},\omega)}{cn(\mathbf{q},\omega)},$ (A-3)
where the relation for the refractive index reads
$n(\mathbf{q},\omega)=\left(\frac{\left|\varepsilon(\mathbf{q},\omega)\right|+\varepsilon_{1}(\mathbf{q},\omega)}{2}\right).$ (A-4)
Figure A.1: Weighted inverse monochromatic opacity [Eq. (A-8)] computed for solid-density hydrogen at $T=4.8~{\rm eV}$ and $\rho=0.08\,{\rm g/cc}$ with momentum transfer $q=0.946\,\textup{~\AA }^{-1}$. The results obtained using the constraints-based noise attenuation method are compared to reference data calculated with $\eta=0.1\,{\rm eV}$, along with results computed for $\eta=0.3\,{\rm eV}$ and $\eta=0.5\,{\rm eV}$. The corresponding DSF data are presented in Fig. 6(a).
@@ A.3 The $\eta$-convergence test using the Kubo-Greenwood dynamic conductivity
The Kubo-Greenwood method is commonly employed to characterize the dynamic transport properties of materials under extreme conditions in the optical limit 114, 74, 115, 97, 98. Given the importance of this approach for such applications as warm dense matter and dense plasmas, we outline the steps to implement the $\eta$-convergence test in the imaginary time domain alongside a constraints-based noise attenuation technique starting from the Kubo-Greenwood conductivity $\sigma_{\rm KG}(\omega)$. In this context, we consider the averaged value of $\sigma_{\rm KG}(\omega)$ defined as the trace of the conductivity tensor 97.
@@
$\chi^{*}_{\rm KG}(\omega)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)\chi_{\rm KG}(\mathbf{q},\omega)\equiv{\rm Im}\left[\frac{1}{1+\frac{4\pi i}{\omega}\sigma_{\rm KG}(\omega)}\right],$ (A-5)
where we introduced the dynamic density response function scaled by $v(q)={4\pi}/{q^{2}}$. We note that $\chi^{*}_{\rm KG}(\omega)$ does not depend on $\mathbf{q}$ because the electron density response exhibits a quadratic dependence $\propto q^{2}$ in the small wavenumber limit for a given fixed configuration of ions 116, 117, 118, 119.
From Eq. (A-5) and Eq. (8), we find the scaled DSF in the Kubo-Greenwood approximation:
$S^{*}_{\rm KG}(\omega)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)S_{\rm KG}(\mathbf{q},\omega)\equiv-\frac{\hbar^{2}}{n}~\frac{1}{1-e^{-\hbar\omega/k_{B}T}}~{\rm Im}\left[\chi^{*}_{\rm KG}(\omega)\right].$ (A-6)
@@
$F^{*}_{\rm KG}(\tau)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)F_{\rm KG}(\tau)\equiv\int_{-\infty}^{\infty}\textnormal{d}\omega\ S^{*}_{\rm KG}(\omega)\ e^{-\tau\omega}.$ (A-7)
@@
After identifying the $\eta$-convergence range, noise filtering can be applied to $S^{*}_{\rm KG}(\omega)$. Following this, the noise-filtered imaginary part of the scaled density response function $\chi^{*}_{\rm KG}(\omega)$ can be computed. The real part of the noise-filtered $\chi^{*}_{\rm KG}(\omega)$ is derived using the Kramers-Kronig relations.
@@
Finally, the noise-attenuated data for the dynamic dielectric function and conductivity in the Kubo-Greenwood approximation can be obtained using the relations $\varepsilon_{\rm KG}^{-1}(\omega)=1+\chi^{*}_{\rm KG}(\omega)$ and $\sigma_{\rm KG}(\omega)=\frac{i\omega}{4\pi}\left[1-\varepsilon_{\rm KG}(\omega)\right]$.
A.4 Monochromatic opacity in solid density hydrogen
As an additional example of the utility of the presented workflow, we consider the weighted monochromatic opacity used in calculating the Rosseland mean opacity. The latter describes radiative transfer due to temperature gradients 120 and is a highly relevant quantity for applications such as stellar evolution models 121. The Rosseland mean opacity $\kappa_{R}$ is a harmonic mean of the frequency-resolved (monochromatic) opacity, weighted by the temperature derivative of the Planck function. It is expressed using the dimensionless variable $u=\omega/T$ as follows 122:
$\frac{1}{\kappa_{R}}=\int_{0}^{\infty}\frac{I(u)}{\kappa(u)}\mathrm{d}u,$ (A-8)
where $I(u)=[15/(4\pi^{4})]u^{4}\exp(-u)/[1-\exp(-u)]^{2}$ and $\kappa(u)=\alpha(u)/\rho$ is the mass absorption coefficient.
In Fig. A.1, we show the results (in atomic units) for the weighted inverse monochromatic opacity $I(u)/\kappa(u)$, which appears in integral (A-8), computed for solid density hydrogen at $T=4.8~{\rm eV}$ and scattering (observation) angle $\theta$ corresponding to the momentum transfer $q=0.946~\textup{~\AA }^{-1}$ via the relation $q=(2\omega_{0}/c)\sin(\theta/2)$ (with $\omega_{0}$ being defined by the XFEL photon energy $E_{0}=\hbar\omega_{0}$). For more information on the conditions for scattering photon energy, please refer to Appendix B of Ref. 50. From Fig. A.1, we observe that applying $\eta$ values outside the convergence range determined by the data analysis in the imaginary time domain results in significant inaccuracies near the maximum. Compared to the reference data with the constraints-based noise attenuation, these inaccuracies lead to errors of $4.6\%$ and $9\%$ in the Rosseland mean opacity $\kappa_{R}$ (A-8) computed using $\eta=0.3~{\rm eV}$ and $\eta=0.5~{\rm eV}$, respectively. This observation highlights the relevance of the presented $\eta$-convergence metric along with the constraints-based noise reduction approach for a broad range of observables beyond the DSF.
@@ A.4 Monochromatic opacity in solid density hydrogen
Figure A.2: Power spectral density of the DSF of aluminum at $T=25~\text{eV}$ for (a) $q=1.0~\textup{~\AA }^{-1}$ and (b) $q=7.75~\textup{~\AA }^{-1}$. The reference data with $\eta=0.5~\text{eV}$ show noticeable noise at high Fourier frequencies, which is significantly reduced through the application of constraint-based filtering.
A.5 The SG filter parameters used for isochorically heated Al
@@
For the DSF of isochorically heated Al ($T=25~{\rm eV}$) at $q=1.0~\textup{~\AA }^{-1}$, we used the third-order polynomial based SG filter with a $w=3.65~{\rm eV}$ window for the data at $\omega\leq 50~{\rm eV}$. At $\omega>50~{\rm eV}$, we used the mean of the results generated by applying the window parameters $w=3.75~{\rm eV}$ and $w=3.5~{\rm eV}$. At $q=7.75~\textup{~\AA }^{-1}$, for the DSF values at $\omega>50~{\rm eV}$, we used the first-order polynomial-based SG filter with the window $w=2.25~{\rm eV}$. For the used filter parameters, we have an efficient high-frequency noise attenuation (as shown in Fig. A.2) and a negligible impact on the ITCF, with both $\bar{F}_{n}(q)/\bar{F}(q,\beta/2)$ and $\sigma_{n}(q)/\bar{F}(q,\beta/2)$ being $10^{-3}[\%]$.
Acknowledgments
@@
Associated Content
@@
The data supporting the findings of this study are available on the Rossendorf Data Repository (RODARE) 123.

### v3

@@ 5 Discussion
State-of-the-art XFEL facilities and modern backlighter set-ups enable the investigation of material properties under new, hitherto unexplored conditions Kraus_Nature_2025, Descamps_sciadv, White_Nature_2025, Tilo_Nature_2023. While advancements in experimental capabilities have greatly enhanced our understanding of high-pressure physics and laser-matter interaction, they also present new challenges for ab initio simulation techniques. It has become crucial to improve the efficiency of ab initio modeling due to the significant increase in computing power that is required to accurately describe the properties of materials under extreme conditions Dornheim_NatureCom_2025, Kononov2023, Ding_prl_2018, White_prl_2020. One key property measured directly in experiments is the XRTS spectrum. Modeling these measurements is a computationally highly intensive task due to strong thermal excitations, the peculiarities of the detectors used to capture the XRTS spectrum, and the variations in temperature and density within the experimental targets. Consequently, developing new computational strategies to improve the efficiency of XRTS modeling has become a crucial task for providing sustainable theoretical support for XRTS experiments, as well as for extending the benchmark and refinement of simplified models commonly used in areas such as radiation hydrodynamics.
In this work, we developed an approach based on the one-to-one mapping between the electronic DSF $S(\mathbf{q},\omega)$ and the ITCF $F(\mathbf{q},\tau)$. This method combines the $\eta$-convergence test in the imaginary time domain with a constraints-based attenuation of narrow-band fluctuations. This combined approach constitutes a highly efficient strategy for XRTS modeling, providing a speed-up of up to an order of magnitude and thus allowing for a significant reduction in the computational cost associated with the analysis of XRTS measurements for targets under extreme conditions. While convergence of the ITCF alone is not sufficient to uniquely determine the DSF, it does ensure the convergence of a number of physical characteristics Dornheim_mre_2023, TC_jcp_2025, including the frequency moments of the DSF Dornheim_prb_2023, tolias2025exactseriesexpansionfrequency, thereby providing a physically grounded basis for the computational strategy presented in this work.
For any given k-point grid, there exists a maximum attainable spectral resolution due to the limited amount of information about dynamic response properties contained in any finite k-point sampling, in contrast to the true thermodynamic limit with an infinite number of states. Besides this limitation, finite k-point sampling introduces narrow-band fluctuations. These fluctuations do not represent physical information but rather constitute an artifact whose amplitude increases as the broadening parameter is reduced. The presented approach improves the fidelity of the DSF by attenuating these identifiable finite-k-point-grid artifacts in a controlled manner, with a significantly smaller impact on resolved physical spectral features than increasing the broadening parameter, while preserving physical characteristics that follow from the ITCF. These benefits are achieved while substantially reducing the required computational cost.
@@
The application of the presented workflow was demonstrated on the example of solid-density hydrogen with partially degenerate electrons under conditions relevant for IFE and laser-driven hydrogen jet experiments. A second example is given by isochorically heated aluminum with hot electrons within an fcc lattice structure of ions. This scenario is pertinent to the physics of transient states studied in experiments that utilize X-ray-driven heating Descamps_sciadv, Descamps2020, Matthieu2011, White2014, Sperling_PRL_2015 (also see discussions in Ref. Moldabekov_prr_2024).
@@
By significantly reducing the computational burden associated with modeling XRTS experiments, the presented approach is expected to play a vital role in computational efforts aimed at supporting new experimental initiatives. In addition to XRTS modeling and interpretation, the presented workflow also enhances the efficiency of modeling a plethora of dynamic response properties that are related to the DSF through the fluctuation-dissipation theorem and the Kramers-Kronig relations. Instead of the DSF, the workflows can be initiated using any dynamic response property that can be expressed in terms of the dynamic density response function  Schoerner_PRE_2023, Ramakrishna_2020, Hamann_PRB_2020, such as dynamic conductivity from the Kubo-Greenwood method Schoerner_PRE_2023, Melton_POP_2024, CALDERIN2017118, PhysRevE.109.065304 (as outlined in Sec. A.3) or TDDFT ullrich2012time, Alex_prl_2025.
@@ A.1 TDDFT simulation details
The TDDFT simulations were carried out using Quantum ESPRESSO Giannozzi_2009, Giannozzi_2017, Giannozzi_jcp_2020, Carnimeo_JCTC_2023, TIMROV2015460, using the Liouville-Lanczos (LL) method to compute the DSF. A static (adiabatic) GGA approximation for the exchange-correlation (XC) kernel was employed in all calculations. For hydrogen, we utilized the pseudopotential h_pbe_v1.4.uspp.F.UPF Garrity_CMS_2014 in conjunction with the PBE XC functional PhysRevLett.77.3865. We considered a density value of $\rho=0.08~\text{g/cc}$ and a temperature of $T=4.8~\text{eV}$, corresponding to the Fermi temperature of the electrons quantum_theory. The size of the simulation box, which contains 14 protons, is set to $L=6.641174\textup{~\AA }$. This temperature leads to a partially degenerate state, with partial ionization Bohme_PRL_2022, Bellenbaum_prr_2025, Hamann_PRR_2023, Bonitz_pop_2024. The DSF results were averaged over 20 snapshots, using a Monkhorst-Pack $k$-point grid set to $10\times 10\times 10$. The convergence of the results with respect to the $k$-point grid was verified by testing an individual snapshot with different $k$-point grids: $8\times 8\times 8$, $10\times 10\times 10$, $14\times 14\times 14$, and $18\times 18\times 18$. The snapshots of protons were randomly selected from the particle trajectories generated by the KSDFT-based molecular dynamics simulations. To obtain $10^{5}$ Lanczos coefficients, we employed a bi-constant extrapolation scheme following the computed number of Lanczos iterations up to $N_{\rm iter}=2\times 10^{4}$ TIMROV2015460. The calculations were executed with a cutoff energy of $75~\text{Ry}$ and $280$ KS orbitals.
For Al, we computed the DSFs in the [100] crystallographic direction and employed the GGA-level PBE functional along with the pseudopotential Al.pbe-sp-van.UPF PhysRevB.77.172102, PP_ESPRESSO_2022, which accounts for 11 valence electrons, and an energy cutoff of 200 Ry. We employed a cubic simulation box containing the conventional fcc unit cell (4 atoms) with an fcc lattice parameter of $a=4.05\,\textup{~\AA }^{-1}$ wyckoff1948crystal. At the temperature of $T=0.025\,\text{eV}$, we used a $40\times 40\times 40$ k-point grid. At the higher temperature of $T=25\,\text{eV}$, we employed two different k-point grids: $10\times 10\times 10$ and $18\times 18\times 18$. Additionally, at $T=25\,\text{eV}$, we utilized 400 KS bands for the cell with 4 Al atoms.
@@ A.2 Connection between different dynamic response functions
Here, we briefly outline the relationships between the DSF and the dynamic conductivity as well as the dynamic dielectric function, which are used to compute the results presented in Section 4.2. For more details, see, e.g., Refs.ginzburg1970propagation, alexandrov1984principles. The DSF is connected to the imaginary part of the density response function through the fluctuation-dissipation theorem, as indicated in Eq. (8). This means that the fluctuation-dissipation theorem and the Kramers-Kronig relations allow one to compute the density response function $\chi(\mathbf{q},\omega)$ from the DSF. The dynamic dielectric function can be expressed in terms of $\chi(\mathbf{q},\omega)$ as
@@ A.3 The $\eta$-convergence test using the Kubo-Greenwood dynamic conductivity
The Kubo-Greenwood method is commonly employed to characterize the dynamic transport properties of materials under extreme conditions in the optical limit Hu_pop_2025, Melton_POP_2024, Maximilian_2023, CALDERIN2017118, PhysRevE.109.065304. Given the importance of this approach for such applications as warm dense matter and dense plasmas, we outline the steps to implement the $\eta$-convergence test in the imaginary time domain alongside a constraints-based filtering starting from the Kubo-Greenwood conductivity $\sigma_{\rm KG}(\omega)$. In this context, we consider the averaged value of $\sigma_{\rm KG}(\omega)$ defined as the trace of the conductivity tensor CALDERIN2017118.
@@
$\chi^{*}_{\rm KG}(\omega)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)\chi_{\rm KG}(\mathbf{q},\omega)\equiv{\rm Im}\left[\frac{1}{1+\frac{4\pi i}{\omega}\sigma_{\rm KG}(\omega)}\right],$ (A-3)
where we introduced the dynamic density response function scaled by $v(q)={4\pi}/{q^{2}}$. We note that $\chi^{*}_{\rm KG}(\omega)$ does not depend on $\mathbf{q}$ because the electron density response exhibits a quadratic dependence $\propto q^{2}$ in the small wavenumber limit for a given fixed configuration of ions Arista_pra_1984, Moldabekov_pop_2018, Moldabekov_Electronic_Structure_2025, Moldabekov_prb_2023.
From Eq. (A-3) and Eq. (8), we find the scaled DSF in the Kubo-Greenwood approximation:
$S^{*}_{\rm KG}(\omega)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)S_{\rm KG}(\mathbf{q},\omega)\equiv-\frac{\hbar^{2}}{n}~\frac{1}{1-e^{-\hbar\omega/k_{B}T}}~{\rm Im}\left[\chi^{*}_{\rm KG}(\omega)\right].$ (A-4)
@@
$F^{*}_{\rm KG}(\tau)\underset{\mathbf{q}\to 0}{=\joinrel=}v(q)F_{\rm KG}(\tau)\equiv\int_{-\infty}^{\infty}\textnormal{d}\omega\ S^{*}_{\rm KG}(\omega)\ e^{-\tau\omega}.$ (A-5)
@@
After identifying the $\eta$-convergence range, filtering of narrow-band fluctuations can be applied to $S^{*}_{\rm KG}(\omega)$. Following this, the imaginary part of the scaled density response function $\chi^{*}_{\rm KG}(\omega)$ filtered from narrow-band fluctuations can be computed. The real part of the response function $\chi^{*}_{\rm KG}(\omega)$ filtered from narrow-band fluctuations is derived using the Kramers-Kronig relations.
@@
Finally, the data with attenuated narrow-band fluctuations for the dynamic dielectric function and conductivity in the Kubo-Greenwood approximation can be obtained using the relations $\varepsilon_{\rm KG}^{-1}(\omega)=1+\chi^{*}_{\rm KG}(\omega)$ and $\sigma_{\rm KG}(\omega)=\frac{i\omega}{4\pi}\left[1-\varepsilon_{\rm KG}(\omega)\right]$.
@@
Figure A.1: Power spectral density of the DSF of aluminum at $T=25~\text{eV}$ for (a) $q=1.0~\textup{~\AA }^{-1}$ and (b) $q=7.75~\textup{~\AA }^{-1}$. The reference data with $\eta=0.5~\text{eV}$ show noticeable narrow-band fluctuations at high Fourier frequencies, which are significantly reduced through the application of constraint-based filtering.
A.4 The SG filter parameters used for isochorically heated Al
@@
For the DSF of isochorically heated Al ($T=25~{\rm eV}$) at $q=1.0~\textup{~\AA }^{-1}$, we used the third-order polynomial based SG filter with a $w=3.65~{\rm eV}$ window for the data at $\omega\leq 50~{\rm eV}$. At $\omega>50~{\rm eV}$, we used the mean of the results generated by applying the window parameters $w=3.75~{\rm eV}$ and $w=3.5~{\rm eV}$. At $q=7.75~\textup{~\AA }^{-1}$ with $10\times 10\times 10$ k-point grid, for the DSF values at $\omega>50~{\rm eV}$, we used the first-order polynomial-based SG filter with the window $w=8.55~{\rm eV}$. For the used filter parameters, we have an efficient attenuation of the narrow-band fluctuations (as shown in Fig. A.1 for the data with $10\times 10\times 10$ k-point grid) and a negligible impact on the ITCF, with both $\bar{F}_{n}(q)/\bar{F}(q,\beta/2)$ and $\sigma_{n}(q)/\bar{F}(q,\beta/2)$ being $10^{-3}[\%]$. For the data computed using $18\times 18\times 18$ k-point grid, we used the first-order polynomial-based SG filter with $w=4.05~{\rm eV}$ at frequencies outside of the range corresponding to 2s-2p feature.
Funding
@@
Competing interest
The authors have no competing interests that might be perceived to influence the results and/or discussion reported in this paper.
Author Contributions
ZM: Conceptualization (lead); Data curation (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Validation (lead); Visualization (lead); Writing – original draft (lead). TD: Conceptualization (equal); Data curation (equal); Formal analysis (equal); Funding acquisition (lead); Methodology (equal); Validation (equal); Writing – original draft (equal). SS, UHA, JV, and MP contributed to data generation, manuscript drafting, and critical discussions. All authors reviewed the manuscript.
@@
The data supporting the findings of this study are available on the Rossendorf Data Repository (RODARE) moldabekov_zhandos_2026_4513.